# Fine-Tuning Multilingual Language Models for Code Review: An Empirical Study on Industrial C# Projects

### v1

@@ 3.6. Evaluation Framework
For Efficiency, we measured the total time taken from receiving the input to producing the output. Human reviewers were timed manually, whereas the response times for LMs and SonarQube were extracted from system log files.
Dataset and Ground Truth. For human-aligned comparisons, we used 40 PRs sampled from Lovion GmbH’s industrial C# repositories. Each PR included one code hunk to ensure alignment. Two senior engineers annotated whether a review comment was required and the associated issue category. Disagreements were resolved through discussion.
Evaluation Procedure. For the 40 PRs:
  • In the Code Change Quality Estimation task, we compared LMs and SonarQube predictions against ground truth and human reviewer judgments (via majority vote).
  • In the Review Comment Generation task, we evaluated the 20 PRs labeled as requiring comments using BLEU and human scores, and further classified each comment as “good” or “not good” based on issue alignment.
Human Reviewer Baseline. We recruited six bilingual software engineers (2 senior, 2 mid-level, 2 junior) to (i) write review comments and (ii) rate model outputs. Their aggregated ratings served as a human baseline across all evaluation dimensions.
@@ 4.1. Results of Code Change Quality Estimation
Figure 4. Code change quality estimation performance of CodeReviewer fine-tuned on monolingual C# data compared to its base multilingual version.
[…]
Bar chart comparing CodeReviewer’s precision, recall, F1, and accuracy for monolingual and multilingual training settings on the code change quality estimation task.
@@ 4.2. Results of Review Comment Generation
DeepSeek-R1-Distill-Llama-8B presents a different trade-off. Fine-tuning on monolingual data significantly boosts BLEU, but slightly reduces Information and Relevance compared to its base model. Notably, the base model achieves the highest Relevance score overall, despite its low BLEU, suggesting that fine-tuning may improve lexical fidelity at the expense of general reasoning and feedback richness.
@@ 4.3. Results of Code Refinement
[…]
Bar chart showing performance comparison of CodeReviewer on code refinement task.
Figure 5. Code refinement performance of CodeReviewer fine-tuned on monolingual C# data compared to its base multilingual version.
@@ 4.4. Comparison of LMs, SonarQube, and Human Reviewers
Figures 6 and 7 present the comparative performance of human reviewers, SonarQube, and LMs across the two main evaluation tasks—code change quality estimation and review comment generation—based on a subset of 40 PRs from our industrial C# dataset.
@@
Review Comment Generation. In the review comment generation task, the performance gap between human reviewers and automated methods became more pronounced (see Figure 7). Humans consistently produced comments with higher Information and Relevance scores across all PRs. Among the LMs, CodeLlama-7B (fine-tuned on C#) achieved the best human evaluation scores, outperforming both SonarQube and the other LMs. Notably, the DeepSeek-R1-Distill-Llama-8B model produced almost identical comments before and after fine-tuning. This suggests that fine-tuning reduced its instruction-following and reasoning capabilities. While its BLEU score increased post-fine-tuning, this improvement came at the cost of reduced explanatory richness and contextual reasoning.
@@
[…]
Comparison of methods on the code change quality estimation task.
Figure 6. Code change quality estimation performance of SonarQube, CodeReviewer and human reviewers on 40 PRs. The CodeReviewer was fine-tuned on C# with English comments.
[…]
Comparison of methods on the comment generation task.
Figure 7. Review comment generation performance of SonarQube, CodeReviewer and human reviewers on 40 PRs. The CodeReviewer was fine-tuned on C# with English comments.
@@ 5. Discussion
5.1. Effectiveness of Monolingual Fine-Tuning for CodeReviewer on Code Change Estimation and Refinement Tasks
@@
These observations highlight a task-dependent trade-off: Monolingual fine-tuning improves classification and structured decision tasks but may hinder performance on semantically rich, generation-oriented tasks.
@@ 5.3. Comparison of Fine-Tuned LMs, SonarQube, and Human Reviewers
Our findings reveal that while monolingual fine-tuning improves LLM performance across code review tasks, human reviewers still consistently outperform both LLM-based models and ASATs like SonarQube. This performance gap was especially pronounced in the Review Comment Generation task, where human comments were rated higher in both Information and Relevance. These results align with prior studies showing that human reviewers provide more actionable, context-aware, and nuanced feedback than current automated tools (Hellendoorn et al., 2021; Cihan et al., 2024).
@@ 6. Implications
Another benefit is cost-efficiency. The evaluated tools—CodeReviewer, CodeLlama-7B, and DeepSeek-R1—are open-source and license-free, providing accessible solutions for organizations operating under budget constraints. Moreover, task-specific monolingual fine-tuning emerged as a practical strategy to improve model effectiveness in single-language environments like C#.
@@ 7. Threats to Validity
Internal Validity. One threat is the limited size and language scope of the monolingual C# dataset. The small volume of high-quality code-review pairs may have restricted the models’ exposure to diverse code patterns, especially in the Code Refinement task. Additionally, translating German review comments into English may have introduced semantic noise, affecting both model learning and human evaluation.
Evaluation variability presents another threat. Although six professional software engineers with varying experience levels rated the outputs, subjective bias remains possible. We mitigated this by averaging scores and resolving annotation disagreements through consensus-based discussion.
@@
Construct Validity. Relying on BLEU for comment evaluation is a known threat, as BLEU captures lexical similarity but may miss semantic quality. To mitigate this, we complemented BLEU with human-centered Information and Relevance ratings. Still, evaluating LLM-generated review quality remains challenging. The choice of baselines may influence results. We used strong open-source models (CodeReviewer, CodeLlama-7B, DeepSeek-R1-Distill-Llama-8B), but comparisons with proprietary models like GPT-4o were beyond our computational resources and budget limitations. Also, for practical reasons, we did not perform multi-run evaluations for each setting, meaning variance across runs remains possible for non-deterministic models like DeepSeek-R1-Distill-Llama-8B.
@@ 8. Conclusion and Future Work
Future research may expand to other object-oriented languages to assess cross-language generalizability. Incorporating structural representations like Abstract Syntax Trees could enhance model understanding. Additionally, applying CoT fine-tuning may improve the reasoning transparency of LLM-generated feedback.

### v2

@@ 3.7.3. Efficiency
Efficiency was measured as the average time-to-review per PR, capturing the complete duration from input to output generation. Human reviewers were timed manually, whereas response times for LMs and SonarQube were extracted from system log files.
@@ 4.1. Results of Code Change Quality Estimation
Figure 2. Code change quality estimation performance of CodeReviewer fine-tuned on monolingual C# dataset compared to its base multilingual version.:
@@ 4.2. Results of Review Comment Generation
DeepSeek-R1-Distill-Llama-8B presents a different trade-off. Fine-tuning on monolingual data significantly boosts BLEU-4, but slightly reduces Information and Relevance compared to its base model. Notably, the base model achieves the highest Relevance score overall, despite its low BLEU-4, suggesting that fine-tuning may improve lexical fidelity at the expense of general reasoning and feedback richness.
@@ 4.3. Results of Code Refinement
Figure 3. Code refinement performance of CodeReviewer fine-tuned on monolingual C# dataset compared to its base multilingual version.:
@@ 4.4. Comparison of LMs, SonarQube, and Human Reviewers
Figures 4 and 5 present the comparative performance of human reviewers, SonarQube, and LMs across the two main evaluation tasks—code change quality estimation and review comment generation—based on a subset of 40 PRs from our industrial C# dataset.
@@
Figure 4. Comparison of SonarQube, CodeReviewer, and human reviewers on 40 PRs in the Code Change Quality Estimation task. The CodeReviewer was fine-tuned on C# with English comments.:
Review Comment Generation. In the review comment generation task, the performance gap between human reviewers and automated methods became more pronounced (see Figure 5). Human reviewers consistently produced comments that correctly addressed the underlying code issues, resulting in the highest Issue Correctness Rate. Among the language models, CodeLlama-7B (fine-tuned on C#) achieved the best alignment with the ground truth, followed closely by CodeReviewer and DeepSeek-R1-Distill-Llama-8B. Notably, the DeepSeek model produced almost identical comments before and after fine-tuning, suggesting that the fine-tuning process reduced its reasoning variety while improving lexical precision. Overall, fine-tuned models narrowed—but did not eliminate—the gap to human reviewers in terms of correctness and issue coverage.
Figure 5. Comparison of SonarQube, CodeReviewer, and human reviewers on 40 PRs in the Review Comment Generation task. The CodeReviewer was fine-tuned on C# with English comments. The Issue Correctness Rate represents the proportion of generated comments judged as “correct,” i.e., accurately addressing the corresponding code issue according to the ground truth.:
@@ 5. Discussion
5.1. Effectiveness of Monolingual Fine-Tuning for CodeReviewer on Code Change Estimation
@@
Overall, these findings demonstrate that domain- and language-specific fine-tuning is particularly beneficial for structured classification-oriented review tasks such as determining whether a code change requires human intervention.
@@ 5.2. Influence of Model Type, PL/NL Scope, and Fine-Tuning Design on Review Comment Generation
5.3. Effectiveness of Monolingual Fine-Tuning for CodeReviewer on Code Refinement
In summary, these observations indicate that fine-tuning is not uniformly beneficial across all ACR stages. While it strengthens discriminative capabilities, it can limit generative diversity and robustness in more complex code transformation tasks.
@@ 5.4. Comparison of Fine-Tuned LMs, SonarQube, and Human Reviewers
Our findings reveal that while monolingual fine-tuning improves LLM performance across code review tasks, human reviewers still consistently outperform both LLM-based models and ASATs like SonarQube. This performance gap was especially pronounced in the Review Comment Generation task, where human reviewers achieved substantially higher semantic accuracy than all automated methods. These results align with prior studies showing that human reviewers provide more actionable, context-aware, and nuanced feedback than current automated tools (Hellendoorn et al., 2021; Cihan et al., 2024).
@@ 6. Implications
Another benefit is cost-efficiency. The evaluated LMs are open-source and license-free, providing accessible solutions for organizations operating under budget constraints. Moreover, task-specific monolingual fine-tuning emerged as a practical strategy to improve model effectiveness in single-language environments like C#.
@@ 7. Threats to Validity
Internal Validity. One threat is the limited size and language scope of the monolingual C# dataset. The small volume of high-quality code-review pairs may have restricted the models’ exposure to diverse code patterns, especially in the Code Refinement task. Additionally, translating German review comments into English may have introduced semantic noise, affecting both model learning and human evaluation. Although translation quality was manually validated with high adequacy, fluency, and inter-annotator agreement, minor translation noise may still persist and could have subtly influenced model behavior.
Evaluation variability presents another threat. Although six professional software engineers with varying experience levels rated the outputs, subjective bias remains possible. We mitigated this by averaging scores and resolving annotation disagreements through consensus-based discussion. To further minimize potential evaluator bias, all raters were independent of the dataset construction and translation validation processes, and human evaluation tasks (manual reviewing vs. model-output assessment) were conducted in separate sessions. Despite these precautions, residual subjective variability cannot be entirely eliminated and represents a remaining internal validity limitation.
@@
Construct Validity. Our choice to employ BLEU metrics follows the same evaluation protocol as the original CodeReviewer study (Li et al., 2022a), ensuring direct comparability with their reported C# results. Although more specialized metrics such as CodeBLEU (Ren et al., 2020) or CrystalBLEU (Eghbali and Pradel, 2023) offer better code-sensitive evaluation, applying them solely to our models—without equivalent outputs from Microsoft’s unpublished checkpoints—would have compromised cross-study comparability. Relying on BLEU-based metrics is a known limitation, as these capture lexical similarity but may overlook semantic or structural correctness. To mitigate this, we complemented BLEU-4 (used for comment generation) and BLEU (used for code refinement) with human-centered Information and Relevance ratings. Still, evaluating LLM-generated review quality remains challenging.
Another construct-related limitation concerns the alignment between refinement outputs and functional correctness. While all generated code snippets were syntactically valid and successfully parsed by the compiler, execution-level or unit-test validation could not be performed due to the absence of complete build contexts in the industrial repositories. As a result, BLEU and EM scores capture lexical and structural similarity but do not fully account for functional or semantic correctness of the generated code.
The choice of baselines may influence results. The choice of baselines may influence results. We used strong open-source models, which represent current state-of-the-art (SOTA) approaches specifically developed for ACR. In contrast, general-purpose frontier LLMs such as GPT-4 or GPT-5 are not explicitly optimized for ACR and cannot be fine-tuned or benchmarked under the same controlled conditions. Therefore, our focus remained on models that are technically comparable, reproducible, and directly aligned with the ACR task objectives. Additionally, comparisons with general-purpose frontier LLMs were beyond our budget constraints. For practical reasons, we did not perform multi-run evaluations for each configuration, meaning that variance across runs remains possible for non-deterministic models like DeepSeek-R1-Distill-Llama-8B.
Also, for practical reasons, we did not perform multi-run evaluations for each setting, meaning variance across runs remains possible for non-deterministic models like DeepSeek-R1-Distill-Llama-8B.
@@ 8. Conclusion and Future Work
Future research may expand to other object-oriented languages to assess cross-language generalizability. Incorporating structural representations like Abstract Syntax Trees could enhance model understanding. Additionally, applying CoT fine-tuning may improve the reasoning transparency of LLM-generated feedback. A systematic analysis of hyperparameter sensitivity could further strengthen the robustness and interpretability of future fine-tuning efforts. Future studies could extend this work by incorporating controlled API-based evaluations of general-purpose SOTA LLMs like GPT-5 to examine whether their broader reasoning and linguistic capacities translate into higher review fidelity. Such experiments would help bridge the gap between domain-specific ACR fine-tuning and general-purpose LLM performance, offering a more complete picture of the trade-offs between specialization and generalization in automated code review.
Future studies could also incorporate statistical significance testing across multiple fine-tuning runs and hyperparameter configurations to quantify the robustness of performance differences. Such analyses would help determine whether observed variations stem from model behavior or stochastic effects during optimization.